# Counting Zeros of Complex-Valued Harmonic Functions via Rouché’s Theorem

### v1

@@ Abstract
Rouché’s Theorem is among the most useful results in complex analysis for counting zeros of analytic functions. Rouché’s Theorem also admits a harmonic analogue for counting zeros of complex harmonic functions. This analogue has previously been applied to closed curves of simple forms, such as circles, to count zeros (e.g., [1]). We apply the harmonic analogue of Rouché’s Theorem directly on non-circular critical curves to determine the total number of zeros of a family of complex harmonic functions given by $f(z)=z^{n}+az^{k}+b\overline{z}^{k}-1$, where $n>k\geq 1$ and $a,b>0$. Under explicit inequalities relating $a$ and $b$, we determine that the total number of zeros is either $n$ or $n+2k$ (counted with multiplicity). We also prove the location of the zeros of $f$ is confined to the union of two explicit annuli in the plane: an inner annulus containing $k$ zeros and an outer annulus containing the remainder. This illustrates how non-circular critical curves can serve as contours in harmonic Rouché arguments for counting and localization.
@@ 1 Introduction
With this example in mind, it is natural to wonder how to count the zeros of $f$ for different parameter values.
Many have sought alternatives to the Fundamental Theorem of Algebra to count zeros of similar families of complex harmonic polynomials, including Brooks et al. [2, 3], Brooks and Lee [5], Sandberg [14], and Work [16]. Among the first were Brilleslyper et al. [1], who studied a family similar to $f$, which they denoted $p_{c}$ and was given by
@@
While investigating, Brilleslyper et al. began focusing on a closed loop called the critical curve, which is the set dividing the complex plane into two regions: (1) sense-preserving and (2) sense-reversing. This idea is further discussed in Section 2.
To count the zeros of $p_{c}$, Brilleslyper et al. successfully applied the argument principle to the critical curve of $p_{c}$. While this argument led to a powerful result for $p_{c}$, it relied heavily upon the critical curve of $p_{c}$ always being circular and also provides little information as to the location of the zeros counted.
Subsequent researchers have studied related families [2, 3, 5, 8, 9, 10, 14, 16], but have focused their efforts on families with circular critical curves or used other methods without directly investigating the critical curve. For more information about the mathematical foundation of this field, see [11, 15].
Recently, researchers such as Melman [12] have begun investigating and bounding the location of zeros of complex analytic and harmonic polynomials through various methods.
Turning our attention back to $f_{1}$ and $f_{2}$, we see that the critical curve (in purple) may take many different shapes beyond that of a single circle, as shown in Figure 4 and Figure 4.
@@
Because little is presently known about how to approach families with such irregular critical curves, we focus our main theorems on (1) counting the zeros of $f$ (see Theorem 1.1 and Theorem 1.2), and (2) identifying two annuli which are guaranteed to contain all zeros of $f$ (see Theorem 1.3). [^1] In this paper, we focus specifically on the case where $a\neq b$. The case where $a=b$ has recently been explored by Brooks et al. [4]. These theorems are given as follows:
@@ 2 Background
(Critical Curve) The set of all $z\in\mathbb{C}$ where $|h^{\prime}(z)|=|g^{\prime}(z)|$ will be called the critical curve.
@@
Before introducing the harmonic analogue of Rouché’s Theorem, we state Rouché’s Theorem for Analytic Functions [13].
@@
We now illustrate how Rouché’s Theorem for Analytic Functions can be used to count the zeros of a complex analytic polynomial in Example 2.4.
Example 2.4.
The analytic polynomial
$\phi_{1}(z)=z^{5}+3z^{2}+1$
has five zeros inside the circle defined by $|z|=2$.
Proof.
Let $p(z)=z^{5}$ and $q(z)=3z^{2}+1$ so that $\phi_{1}=p+q$. Take $C$ to be the closed contour $|z|=2$. Then, for $z$ on $C$,
$|p(z)|=|z^{5}|=|z|^{5}=2^{5}=32.$
Similarly, using the triangle inequality,
$|q(z)|$ $\leq|3z^{2}|+1$
$=3|z|^{2}+1$
$=3(2)^{2}+1$
$=13.$
Since $p$ and $q$ are analytic in the open set containing the simple closed contour $C$ and its interior, and $|p(z)|>|q(z)|$ for all z on $C$, $p$ and $p+q$ must have the same number of zeros inside $C$. Because $p$ has five zeros located within $C$, $\phi_{1}$ also has five zeros in $C$.
Figure 5: Zeros of $\phi_{1}$ within $C$
We confirm the proof of Example 2.4 graphically by plotting the five zeros of $\phi_{1}$ found within $C$ in Figure 5.
Similar to the previous example, we may count the zeros of $z^{5}+\overline{z}^{3}-1$ by applying the harmonic analogue of Rouché’s Theorem.
@@
This theorem is proved similarly to the proof of Rouché’s Theorem by Brown and Churchill [6].
We see Rouché’s Theorem is a powerful tool in counting zeros within a closed curve. Notably, Rouché’s Theorem has historically been used with primarily simple closed loops, such as the circle illustrated in Figure 5.
Returning to $f$, it is not immediately clear how Rouché’s Theorem helps count zeros. However, if Rouché’s Theorem could be applied to determine (1) the sum of the orders of the zeros of $f$, as well as (2) how many zeros appear within the sense-reversing region of $f$ (i.e., within its critical curve), the total number of zeros of $f$ could quickly be deduced.
The first proof to determine the sum of the orders of zeros of $f$ is relatively straightforward and is proved as Lemma 3.1 in Section 3.
The second proof is more complicated. Consider once again Figure 4 and Figure 4, given in Section 1. We see the critical curve of $f$ may take different shapes and may have one or more closed loops. This indicates applying a traditional Rouché’s Theorem argument directly to the critical curve with a constant value of $|z|$ is insufficient.
The first pair of theorems in Section 3 are dedicated to counting the zeros of $f$ by applying Rouché’s Theorem for Harmonic Functions directly to the critical curve of $f$. Our final theorem deduces two annuli which contain all zeros of $f$.
@@ 3.1 Zero Counting
$|h(z)|$ $<|z|^{k}\left(\frac{bk(1+\lambda)}{n}+b\lambda+\epsilon\right)$
@@ 4 Conclusion
We have shown how Rouché’s Theorem can be applied to non-circular critical curves to determine the exact number of zeros of a complex harmonic polynomial. It is fair to wonder how many more types of critical curves can be analyzed in this way. Our current analysis generalizes similarly whenever $n$ is a natural number with $n>k$, including when $k$ is negative (i.e., when $f$ has poles).
In the future, interesting results would be to determine exact values for $b_{0}$ such that different $\lambda$ values determine the exact number of zeros of $f$, beyond the asymptotic behavior of $f$ as $a$ or $b$ increase toward infinity. This may require additional analysis techniques.
It is also believed that stronger bounds can be formed to locate the zeros, specifically that the bound of $R_{2}$ in Theorem 1.3 can be sharpened to $R_{2}=\frac{1}{(b+a-1)^{\frac{1}{k}}}$. Moreover, it is suspected that sector bounds could also be found to further reduce the area in the plane where zeros of $f$ might be located.
@@ Acknowledgements
I would like to thank my research advisor, Dr. Jennifer Brooks, for going above and beyond in her feedback, guidance and mentorship. I would also like to thank former group members, Alex Lee, Chammelia Zentz and George Westover, for fruitful discussions during the early stages of forming the results presented in this paper.

### v2

@@ Abstract
Rouché’s Theorem is among the most useful results in complex analysis for counting zeros of analytic functions. Rouché’s Theorem also admits a harmonic analogue for counting zeros of complex harmonic functions. Previously, this analogue has been applied primarily to closed curves of simple geometry, such as circles, to count zeros. We demonstrate that non-circular critical curves can serve as effective contours by applying a harmonic Rouché-type argument to determine the total number of zeros of the complex harmonic family given by $f(z)=z^{n}+az^{k}+b\overline{z}^{k}-1$, where $n>k\geq 1$ and $a,b>0$. Under explicit inequalities relating $a$ and $b$, we determine the total number of zeros is either $n$ or $n+2k$ (counted with multiplicity). We also prove the zeros of $f$ are confined to the union of two explicit annuli in the plane: an inner annulus containing $k$ zeros and an outer annulus containing the remainder.
@@ 1 Introduction
With this example in mind, it is natural to wonder how to count the zeros of $f$ for different parameter values, as well as what the possible number of zeros of $f$ could be.
While investigating complex harmonic polynomials of degree $n$, Wilmshurst [15] conjectured the valence, and therefore the count of the zeros, for such polynomials to be bounded above by $n^{2}$. At the same time, Wilmshurst constructed examples achieving this bound, indicating it is sharp. However, Wilmshurst’s conjecture did not yet provide the tools to find the exact count of zeros or their location for specific harmonic polynomial families.
Among the first to count the zeros of a specific harmonic family were Brilleslyper et al. [1], who studied the harmonic trinomial family, denoted $p_{c}$, given by
@@
In their analysis, Brilleslyper et al. focused on a closed loop called the critical curve, which is the set dividing the complex plane into two regions: (1) sense-preserving and (2) sense-reversing. This idea will be discussed further in Section 2.
To count the zeros of $p_{c}$, Brilleslyper et al. successfully applied the argument principle to the critical curve of $p_{c}$. While this argument led to a precise zero-counting result for their family, it relied heavily upon the geometry of the critical curve of $p_{c}$ being circular and also provided little information as to the location of the zeros counted.
Subsequent researchers have studied harmonic trinomial families [2, 3, 16], including related families with poles [5, 10], but have similarly focused their efforts on families with circular critical curves or used other methods without directly analyzing the critical curve. Sandberg [14] recently investigated a specific family of harmonic quadrinomials constructed so that the unit circle is always at least part of the critical curve, providing both the count of the zeros and bounds for their locations for given parameter values. Even so, results beyond harmonic trinomials and harmonic families with circular critical curves remained limited. For more information about the mathematical foundation of this field, see [11, 15].
Also of interest, researchers have begun investigating and bounding the location of zeros of complex harmonic polynomials through various methods. Recently, Geleta and Alemu [8] derived the locations for the zeros of $p_{c}$ and provided a closed disk in which all zeros of complex-valued harmonic polynomials must exist. More detailed regions for harmonic trinomials were later found by Gao et al. [9] and Melman [12], opening the door to similarly investigate complex harmonic quadrinomials with more complicated geometries.
Turning our attention back to our earlier example, we see the quadrinomial
$f(z)=z^{n}+az^{k}+b\overline{z}^{k}-1$
is constructed by introducing a fourth term, $az^{k}$, to the family $p_{c}$ studied by Brilleslyper et al. Notably, the introduction of this term provides the simplest harmonic extension of $p_{c}$ that breaks circular symmetry of the critical curve, as evidenced in Figure 4 and Figure 4.
More specifically, by separating $f$ into its analytic ($h$) and co-analytic ($\overline{g}$) parts, the critical curve of $f$ is defined by the set of all points in $\mathbb{C}$ where $|h^{\prime}(z)|=|g^{\prime}(z)|$, i.e., $|nz^{n-1}+akz^{k-1}|=kb|z|^{k-1}$, which forms weighted lemniscate-type curves similar to those previously analyzed by Melman [11] in both analytic and harmonic settings. Similar geometries arise where multiple analytic and co-analytic monomials interact. Thus, introducing a second analytic monomial term in $f$ provides a natural minimal test case for determining whether the harmonic analogue of Rouché’s Theorem may be used to extend zero-counting techniques beyond families of harmonic trinomials (or those with otherwise circular geometries) to more general harmonic polynomials.
@@
Because the behavior of harmonic polynomials with non-circular critical curves remains poorly understood, we address this gap by analyzing the complex-valued harmonic quadrinomial $f$. This paper contributes a generalization to count the zeros of a harmonic polynomial without circular symmetry and provides two annuli guaranteed to contain all zeros of $f$. We therefore focus our main theorems on (1) counting the zeros of $f$ (see Theorem 1.1 and Theorem 1.2), and (2) identifying two annuli which must contain all zeros of $f$ (see Theorem 1.3). [^1] In this paper, we focus specifically on the case where $a\neq b$. The case where $a=b$ has recently been explored by Brooks et al. [4]. These theorems are given as follows:
@@ 2 Background
(Critical Curve) The set of all $z\in\mathbb{C}$ where $|h^{\prime}(z)|=|g^{\prime}(z)|$ will be called the critical curve of $f$.
@@
We first recall Rouché’s Theorem for Analytic Functions [13].
@@
The harmonic analogue of Rouché’s Theorem counts zeros with respect to their order and multiplicity, and is proved using a similar approach to the proof in the analytic setting by Brown and Churchill [6].
@@
We see Rouché’s Theorem is a powerful tool for counting zeros within a closed curve. Notably, while Rouché’s Theorem has typically been applied to simple closed loops such as circles, it may be used on any simple closed curve in $\mathbb{C}$ on which its hypotheses are satisfied.
Returning to $f$, it is not immediately clear how Rouché’s Theorem can be used to count its zeros. However, if Rouché’s Theorem can be applied to determine (1) the sum of the orders of the zeros of $f$, as well as (2) how many zeros lie within the sense-reversing region of $f$ (i.e., within its critical curve), then the total number of zeros of $f$ follows quickly. The first of these quantities is straightforward to obtain and is proved as Lemma 3.1 in Section 3.
The second quantity requires more attention. Consider once again Figure 4 and Figure 4, given in Section 1. We see the critical curve of $f$ may take different shapes and may have one or more closed loops. This shows that applying a traditional Rouché argument on circles of constant modulus is insufficient to count the zeros that lie within the critical curve.
Accordingly, the first pair of theorems in Section 3 is dedicated to counting the zeros of $f$ by applying Rouché’s Theorem for Harmonic Functions directly to the critical curve of $f$. Our final theorem identifies two annuli that contain all zeros of $f$.
@@ 3.1 Zero Counting
$|h(z)|$ $<|z|^{k}\left(\frac{bk(1+\lambda)}{n}+b\lambda+\epsilon\right).$
@@ 4 Conclusion
We have shown how Rouché’s Theorem can be applied to non-circular critical curves to determine the exact number of zeros of a complex harmonic polynomial. It is natural to ask how broadly this approach can be extended to other types of critical curves. Our analysis generalizes similarly whenever $n$ is a natural number with $n>k$, including cases where $k$ is negative (i.e., when $f$ has poles).
In the future, interesting directions include determining exact values for $b_{0}$ for which different $\lambda$-values determine the exact number of zeros of $f$, beyond the asymptotic behavior of $f$ as $a$ or $b$ increase toward infinity. This may require additional analytical techniques.
It is also believed that stronger bounds can be formed to locate the zeros, specifically that the bound of $R_{2}$ in Theorem 1.3 can be sharpened to $R_{2}=\frac{1}{(b+a-1)^{\frac{1}{k}}}$. Moreover, it is suspected that sector bounds could also be found to further reduce the region in the plane where zeros of $f$ might be located.
@@ Acknowledgements
I would like to thank my research advisor, Dr. Jennifer Brooks, for going above and beyond in her feedback, guidance, and mentorship. I would also like to thank former group members, Alex Lee, Chammelia Zentz, and George Westover, for fruitful discussions during the early stages of forming the results presented in this paper.